a stone is dropped from certaimheight which can reach the ground in 5s if the stine is dropped after 3seconds of its fall and then allowed to fall again then the time taken by stone to reach the ground for the remaining distace?

Grade:

1 Answers

Gokul Joshi AskiitiansExpert-IITK
13 years ago

Dear Sri,

I think the correct question would be:
a stone is dropped from certain height and reaches the ground in 5s. If the stone is stopped after 3 seconds of its fall and then allowed to fall again then what is the time taken by stone to reach the ground for the remaining distance?

H = 1/2 g t2 = 1/2 * 10 * 52 = 125 m // total height
H3 = 1/2 g t32 = 1/2 * 10 * 32 = 45 m // drop for 3 secs
remaining = 125 - 45 = 80 m = 1/2 g t02 = > t0 = 4 sec

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly. All the best.

Regards,
Askiitians Experts,
Gokul  Joshi